I need to solve 'a' for this equation:
4a + 38 = 20a - 106

Answer by rothauserc(4718)

4a + 38 = 20a - 106
subtract 4a from both sides of =
38 = 16a - 106
add 106 to both sides of =
16a = 144
a = 144 / 16 = 9
therefore,
a = 9
*****************************************
check the answer
4*9 + 38 = 20*9 - 106
36 + 38 = 180 - 106
74 = 74
our answer checks :-)